# When Federated Learning Meets Quantum Computing: Survey and Research Opportunities

### v2

@@ 7 Privacy-Preserving QFL
[102] Robust post-quantum security with efficient 3-round protocol; lower computational load Limited analysis on dropout handling and large-scale scalability Moderate [optimized homomorphic encryption] Synchronous aggregation
@@
[103] Integrates fully homomorphic encryption, achieving 40× training efficiency boost; strong AUC performance High computational overhead of FHE; deployment challenges in NISQ settings High [matrix-vector encoding] Synchronous aggregation
@@
[104] Introduces controlled quantum noise to enhance adversarial robustness; effective on high-dimensional data Performance degrades beyond optimal noise levels; lacks extensive real-world scaling tests Moderate [noise injection strategy] Asynchronous aggregation
@@
[105] Blind quantum computing ensures strong privacy; robust under gradient attacks High computational cost; dependency on advanced quantum infrastructure Moderate [blind computing methods] Synchronous aggregation
@@
[53] Fully asynchronous updates; strong privacy via Paillier encryption High encryption overhead; limited dropout handling Moderate [Paillier-based aggregation] Asynchronous aggregation
@@
[106] Achieves $>98\%$ test accuracy with $\epsilon<1.3$; strong privacy protection Noise trade-offs impact performance; scaling challenges High [noise-resilient VQC] Asynchronous aggregation
@@
[107] Two distinct protocols with low communication costs; secure gradient protection via quantum states Practical deployment challenges; noise in quantum communication Moderate [efficient state encoding] Synchronous aggregation
@@
[108] Lattice-based scheme reduces computational overhead by 20%, double masking enhances privacy Multi-stage reconstruction complexity; limited scalability for extensive participants Moderate [lattice cryptography] Synchronous aggregation
@@
[109] GHZ-based aggregation ensures privacy; low computational complexity without encryption overhead Dependence on quantum infrastructure; scalability to large datasets not demonstrated High [GHZ state aggregation] Synchronous aggregation
@@
Privacy-preserving mechanisms in QFL are essential for safeguarding sensitive information while leveraging the computational advantages of quantum systems. QFL introduces advanced techniques such as quantum differential privacy [48, 110], secure multi-party computation [111, 108], and homomorphic encryption [59, 103] to mitigate threats like gradient leakage and model inversion. These methodologies not only enhance data security but also address the unique challenges posed by quantum noise, ensuring robust performance in sensitive applications such as healthcare, finance, and autonomous systems.
@@
$\bullet$ Challenges: Implementing privacy-preserving mechanisms in QFL presents unique challenges that intertwine the complexities of QC with stringent data confidentiality requirements. A significant concern is the potential for quantum-specific attacks, such as quantum gradient inversion, where adversaries exploit quantum information to reconstruct sensitive data from shared gradients [112]. Unlike classical systems, the quantum nature of data and computations necessitates novel privacy-preserving protocols that can effectively conceal gradient information without compromising the advantages of quantum processing.
@@
Moreover, the integration of Differential Privacy (DP) into QFL frameworks is not straightforward. The inherent noise in NISQ devices complicates the application of traditional DP techniques [113]. Additionally, the development of quantum-compatible encryption methods, such as quantum homomorphic encryption, remains in its infancy, presenting challenges in performing computations on encrypted data without exposing sensitive information [114]. These limitations underscore the need for advanced research into quantum-specific privacy-preserving techniques that can balance the trade-offs between data security and computational efficiency in FL environments.
@@ 7.1 Quantum Differential Privacy in Federated Systems
Quantum Differential Privacy (QDP) plays a vital role in protecting sensitive data in FL systems, particularly when integrated with quantum computing benefits. A protocol using blind quantum computing has been introduced by [105] for private QFL, which utilizes quantum servers for intensive computations and ensures data privacy through differential privacy. However, a drawback is the possible decline in performance when extending this protocol to more intricate datasets and non-IID data distributions. The work in [106] discussed the challenges of integrating QFL with QDP on NISQ devices and emphasized the importance of balancing privacy preservation with model performance. An innovative quantum fuzzy FL framework is introduced by [60], which amalgamates fuzzy logic with quantum FL to augment both privacy and model efficacy.
@@
Using FHE, the authors in [103] proposed high-performance and quantum-safe federated learning, a vertical FL strategy, to achieve quantum-resistant training. One of the most compelling aspects of the study is the substantial boost in training efficiency, reported as up to 40 times faster than conventional methods while maintaining an Area Under the Curve (AUC) comparable to standard algorithms like logistic regression and XGBoost. Furthermore, the work [104] investigated the use of quantum noise to improve model robustness in FL, particularly against adversarial attacks; and through extensive numerical simulations, it has been demonstrated that the introduction of controlled quantum noise can enhance the resilience of the model.
@@
An assessment of $\mathrm{QUE}$ for privacy‐preserving QFL protocols indicates that frameworks prioritizing encryption‐heavy workflows tend to operate with moderate qubit–layer occupancy ($\mathrm{QUE}\approx 0.6$–0.8), while those with lighter embedding circuits achieve higher occupancy. For instance, the FHE‐driven vertical update in [103] typically runs on a four‐qubit, three‐layer VQC ($\mathrm{Q}_{alloc}=4$, $L=3$), yielding $\mathrm{QUE}\approx 0.67$ and justifying its synchronous aggregation to amortize the cost of homomorphic operations. The differential‐privacy–enhanced protocol in [106] uses a similar shallow ansatz ($\mathrm{Q}_{alloc}=5$, $L=3$), achieving $\mathrm{QUE}\approx 0.67$ and favoring asynchronous updates to hide noise patterns across rounds. Blind computing methods in [105] deploy deeper circuits ($\mathrm{Q}_{alloc}=6$, $L=4$), giving $\mathrm{QUE}\approx 0.75$ and employing synchronous UBQC updates to maintain coherent secret sharing. Noise‐adaptive training in [104] uses a three‐qubit, two‐layer ansatz ($\mathrm{Q}_{alloc}=3$, $L=2$), yielding $\mathrm{QUE}\approx 1.0$ and leveraging asynchronous updates to inject and calibrate noise on the fly. Finally, the Paillier‐secured asynchronous framework in [53] runs on a minimal two‐qubit, two‐layer circuit ($\mathrm{QUE}=1.0$), matching its lock‐free aggregation for low‐latency privacy protection. Despite their strengths, these protocols often lack standardized metrics for end-to-end latency and energy cost, critical metrics for real-world deployment, and none address dynamic adjustment of privacy budgets in response to fluctuating noise levels. Future work should establish unified benchmarking frameworks, investigate adaptive noise schedules tuned to qubit coherence times, and explore lightweight encryption primitives that optimize the trade-off between QUE and cryptographic robustness.
@@ 7.2 Quantum Model Aggregation and Attacks
The importance of privacy-preserving model aggregation has become pronounced with the continuous expansion of FL, particularly within quantum computing contexts. In [53], a concept of privacy-preserving asynchronous FL has been introduced to facilitate secure model aggregation without necessitating the synchronization of client updates. The authors employed Paillier encryption [115] to support the homomorphic aggregation of encrypted model updates. Here, clients can update asynchronously by contributing local models, with the server aggregating these updates without disclosing private information. Addressing the need for post-quantum security, the work [102] presented a secure three-round aggregation protocol to ensure quantum resistance, using post-quantum cryptographic techniques.
Figure 9: Illustrating common attacks on FL and QFL frameworks.
@@
One prominent direction concerns safeguarding FL infrastructures against quantum attacks. Fig. 9 illustrates various attacks on quantum, FL, and QFL sides. The work [108] introduced an FL framework that utilizes lattice-based multi-stage secret sharing to ensure resistance against quantum attacks. This scheme employs dual masking techniques to safeguard model parameters, enabling reconstruction of masked model updates while minimizing frequent communication overhead. On a similar track, the authors in [116] introduced a framework that leverages post-quantum cryptographic primitives to mitigate eavesdropping and model-tampering threats. By combining ratcheting methods with on-chain transactions, such a framework provides forward secrecy for iterative FL updates while recording contributions in an immutable ledger. Parallel to security concerns, researchers have explored adversarial defenses within QFL [117], which integrates adversarial training in a quantum federated environment, generating local perturbations at each node and sharing robust gradients via FedAvg. Using MNIST, they indicated a $12\%$ improvement in accuracy under moderate projected gradient descent attacks when at least half of the clients ran adversarially trained subroutines. This highlights the merit of local adversarial coverage even in small QNNs.
@@
From an attack perspective, to counter model poisoning and gate‐parameter manipulation in QFL, operators can adopt Byzantine‐robust aggregation rules and runtime validation of quantum updates. Classical FL defenses such as Krum or coordinate‐wise median have proven effective against malicious client updates [118], and can be extended to QFL by treating each client’s variational parameter shift as a vector for robust aggregation. Additionally, clients and the server can interleave randomized “test circuits” which would essentially be short sequences with known output distributions—into the training pipeline. By measuring the fidelity of these test circuits against expected outcomes using the average‐gate‐fidelity formula [119], the server can detect aberrant gate manipulations or backdoor injections before integrating a client’s update.
@@
Quantum‐specific error mitigation techniques further strengthen security by masking subtle parameter perturbations. Techniques like zero‐noise extrapolation or probabilistic error cancellation [39] can be applied selectively to aggregated model parameters, reducing the impact of malicious noise injections while preserving the integrity of genuine updates. From a cryptographic standpoint, integrating verifiable blind QC protocols [105] allows the server to challenge clients with hidden circuit specifications, guaranteeing that parameterized unitaries are executed as intended without revealing the model’s secrets. Together, these measures: robust aggregation, fidelity‐based auditing, error‐mitigation filtering, and verifiable execution, offer a layered defense tailored to the hybrid vulnerabilities of QFL.
@@
An assessment of $\mathrm{QUE}$ for secure‐aggregation and attack‐resilient QFL schemes shows uniformly high occupancy due to their use of shallow embedding circuits. For example, the Paillier‐based asynchronous framework in [53] runs on a minimal two‐qubit, two‐layer ansatz ($\mathrm{Q}_{alloc}=2$, $L=2$), yielding $\mathrm{QUE}=1.0$ and matching its lock‐free aggregation strategy. Post‐quantum MPC in [102] and lattice‐masked secret sharing in [108] both employ four‐qubit, three‐layer circuits ($\mathrm{Q}_{alloc}=4$, $L=3$) with $\mathrm{QUE}=1.0$, justifying synchronous updates to amortize cryptographic overhead. The blockchain‐anchored protocol in [116] uses a three‐qubit, two‐layer design ($\mathrm{QUE}=1.0$) and also opts for synchronous aggregation to ensure ledger consistency. The adversarial defense in [117] leverages a four‐qubit, four‐layer QNN ($\mathrm{QUE}=1.0$) and employs asynchronous FedAvg with local perturbation to balance robustness and throughput. Despite their strong QUE and clear strategy alignment, these protocols share key limitations. These models generally ignore client dropout and dynamic participation, lack adaptive privacy‐budget tuning, and incur substantial computational and communication costs from heavy cryptographic primitives. Future work should explore lightweight quantum‐safe primitives, robust dropout‐tolerant aggregation, and real‐time privacy budget adjustment to enhance scalability and resilience in practical QFL deployments.

### v4

@@ 7 Privacy-Preserving QFL
[QFL-SecAggPostQ1] Robust post-quantum security with efficient 3-round protocol; lower computational load Limited analysis on dropout handling and large-scale scalability Moderate [optimized homomorphic encryption] Synchronous aggregation
@@
[QFL-DPFHE] Integrates fully homomorphic encryption, achieving 40× training efficiency boost; strong AUC performance High computational overhead of FHE; deployment challenges in NISQ settings High [matrix-vector encoding] Synchronous aggregation
@@
[QFL-DPNoise] Introduces controlled quantum noise to enhance adversarial robustness; effective on high-dimensional data Performance degrades beyond optimal noise levels; lacks extensive real-world scaling tests Moderate [noise injection strategy] Asynchronous aggregation
@@
[QFL-BlindC] Blind quantum computing ensures strong privacy; robust under gradient attacks High computational cost; dependency on advanced quantum infrastructure Moderate [blind computing methods] Synchronous aggregation
@@
[QFL-AsynchAgg] Fully asynchronous updates; strong privacy via Paillier encryption High encryption overhead; limited dropout handling Moderate [Paillier-based aggregation] Asynchronous aggregation
@@
[QFL-DP] Achieves $98\%$ test accuracy with $\epsilon<1.3$; strong privacy protection Noise trade-offs impact performance; scaling challenges High [noise-resilient VQC] Asynchronous aggregation
@@
[QFL-GradientDP] Two distinct protocols with low communication costs; secure gradient protection via quantum states Practical deployment challenges; noise in quantum communication Moderate [efficient state encoding] Synchronous aggregation
@@
[QFL-SecAggPostQ2] Lattice-based scheme reduces computational overhead by 20%, double masking enhances privacy Multi-stage reconstruction complexity; limited scalability for extensive participants Moderate [lattice cryptography] Synchronous aggregation
@@
[QFL10] GHZ-based aggregation ensures privacy; low computational complexity without encryption overhead Dependence on quantum infrastructure; scalability to large datasets has not been demonstrated High [GHZ state aggregation] Synchronous aggregation
@@
Privacy-preserving mechanisms in QFL are essential for safeguarding sensitive information while leveraging the computational advantages of quantum systems. QFL introduces advanced techniques such as quantum differential privacy [hirche2023quantum, zhou2017differential], secure multi-party computation [crepeau2002secure, QFL-SecAggPostQ2], and homomorphic encryption [broadbent2015quantum, QFL-DPFHE] to mitigate threats like gradient leakage and model inversion. These methodologies not only enhance data security but also address the unique challenges posed by quantum noise, ensuring robust performance in sensitive applications such as healthcare, finance, and autonomous systems.
@@
$\bullet$ Challenges: Implementing privacy-preserving mechanisms in QFL presents unique challenges that intertwine the complexities of QC with stringent data confidentiality requirements. A significant concern is the potential for quantum-specific attacks, such as quantum gradient inversion, where adversaries exploit quantum information to reconstruct sensitive data from shared gradients [larson2020quantum]. Unlike classical systems, the quantum nature of data and computations necessitates novel privacy-preserving protocols that can effectively conceal gradient information without compromising the advantages of quantum processing.
@@
Moreover, the integration of Differential Privacy (DP) into QFL frameworks is not straightforward. The inherent noise in NISQ devices complicates the application of traditional DP techniques [zhao2024bridging]. Additionally, the development of quantum-compatible encryption methods, such as quantum homomorphic encryption, remains in its infancy, presenting challenges in performing computations on encrypted data without exposing sensitive information [fisher2014quantum]. These limitations underscore the need for advanced research into quantum-specific privacy-preserving techniques that can balance the trade-offs between data security and computational efficiency in FL environments.
@@ 7.1 Quantum Differential Privacy in Federated Systems
Quantum Differential Privacy (QDP) plays a vital role in protecting sensitive data in FL systems, particularly when integrated with quantum computing benefits. A protocol using blind quantum computing has been introduced by [QFL-BlindC] for private QFL, which utilizes quantum servers for intensive computations and ensures data privacy through differential privacy. However, a drawback is the possible decline in performance when extending this protocol to more intricate datasets and non-IID data distributions. The work in [QFL-DP] discussed the challenges of integrating QFL with QDP on NISQ devices and emphasized the importance of balancing privacy preservation with model performance. An innovative quantum fuzzy FL framework is introduced by [QFL-QFuzzyNN], which amalgamates fuzzy logic with quantum FL to augment both privacy and model efficacy.
@@
Using FHE, the authors in [QFL-DPFHE] proposed high-performance and quantum-safe federated learning, a vertical FL strategy, to achieve quantum-resistant training. One of the most compelling aspects of the study is the substantial boost in training efficiency, reported as up to 40 times faster than conventional methods while maintaining an Area Under the Curve (AUC) comparable to standard algorithms like logistic regression and XGBoost. Furthermore, the work [QFL-DPNoise] investigated the use of quantum noise to improve model robustness in FL, particularly against adversarial attacks.
@@
An assessment of $\mathrm{QUE}$ for privacy‐preserving QFL protocols indicates that frameworks prioritizing encryption‐heavy workflows tend to operate with moderate qubit–layer occupancy ($\mathrm{QUE}\approx 0.6$–0.8), while those with lighter embedding circuits achieve higher occupancy. For instance, the FHE‐driven vertical update in [QFL-DPFHE] typically runs on a four‐qubit, three‐layer VQC ($\mathrm{Q}_{alloc}=4$, $L=3$), yielding $\mathrm{QUE}\approx 0.67$ and justifying its synchronous aggregation to amortize the cost of homomorphic operations. The differential‐privacy–enhanced protocol in [QFL-DP] uses a similar shallow ansatz ($\mathrm{Q}_{alloc}=5$, $L=3$), achieving $\mathrm{QUE}\approx 0.67$ and favoring asynchronous updates to hide noise patterns across rounds. Blind computing methods in [QFL-BlindC] deploy deeper circuits ($\mathrm{Q}_{alloc}=6$, $L=4$), giving $\mathrm{QUE}\approx 0.75$ and employing synchronous UBQC updates to maintain coherent secret sharing. Noise‐adaptive training in [QFL-DPNoise] uses a three‐qubit, two‐layer ansatz ($\mathrm{Q}_{alloc}=3$, $L=2$), yielding $\mathrm{QUE}\approx 1.0$ and leveraging asynchronous updates to inject and calibrate noise on the fly. Finally, the Paillier‐secured asynchronous framework in [QFL-AsynchAgg] runs on a minimal two‐qubit, two‐layer circuit ($\mathrm{QUE}=1.0$), matching its lock‐free aggregation for low‐latency privacy protection. Despite their strengths, these protocols often lack standardized metrics for end-to-end latency and energy cost critical metrics for real-world deployment.
@@ 7.2 Quantum Model Aggregation and Attacks
The importance of privacy-preserving model aggregation has become pronounced with the continuous expansion of FL, particularly within quantum computing contexts. In [QFL-AsynchAgg], a concept of privacy-preserving asynchronous FL has been introduced to facilitate secure model aggregation without necessitating the synchronization of client updates. The authors employed Paillier encryption [paillier2005paillier] to support the homomorphic aggregation of encrypted model updates. Here, clients can update asynchronously by contributing local models, with the server aggregating these updates without disclosing private information. Addressing the need for post-quantum security, the work [QFL-SecAggPostQ1] presented a secure three-round aggregation protocol to ensure quantum resistance, using post-quantum cryptographic techniques.
Figure 10: Illustrating attacks on FL and QFL frameworks.
@@
One prominent direction concerns safeguarding FL infrastructures against quantum attacks. Fig. 10 illustrates various attacks on quantum, FL, and QFL sides. The work [QFL-SecAggPostQ2] introduced an FL framework that utilizes lattice-based multi-stage secret sharing to ensure resistance against quantum attacks. This scheme employs dual masking techniques to safeguard model parameters, enabling reconstruction of masked model updates while minimizing frequent communication overhead. On a similar track, the authors in [gharavi2025pqbflpostquantumblockchainbasedprotocol] introduced a framework that leverages post-quantum cryptographic primitives to mitigate eavesdropping and model-tampering threats. Parallel to security concerns, researchers have explored adversarial defenses within QFL [maouaki2025qfalquantumfederatedadversarial], which integrates adversarial training in a quantum federated environment, generating local perturbations at each node and sharing robust gradients via FedAvg.
@@
From an attack perspective, to counter model poisoning and gate‐parameter manipulation in QFL, operators can adopt Byzantine‐robust aggregation rules and runtime validation of quantum updates. Classical FL defenses such as Krum or coordinate‐wise median have proven effective against malicious client updates [Blanchard2017Krum], and can be extended to QFL by treating each client’s variational parameter shift as a vector for robust aggregation. Additionally, clients and the server can interleave randomized “test circuits,” which would essentially be short sequences with known output distributions, into the training pipeline. By measuring the fidelity of these test circuits against expected outcomes using the average‐gate‐fidelity formula [QC9], the server can detect aberrant gate manipulations or backdoor injections before integrating a client’s update.
Further, from an aspect of a distributed denial-of-service (DDoS), the attacker can target three pressure points: the central aggregator, Quantum-Compute-as-a-Service (QCaaS) endpoints for PQC evaluations, and intermediate relays that coordinate submissions and returns. Congestion at any of these points reduces effective throughput, inflates staleness $\Delta t_{k}$, and, most critically for privacy, shrinks the instantaneous participant set $\mathcal{S}_{t}$. Secure aggregation delivers its guarantees only when the active set meets a minimum threshold; letting rounds proceed with too few contributors erodes anonymity and exposes client updates.
To make the dependency explicit, let $k_{\min}$ denote the minimum cohort size required by the secure aggregation protocol. A round is privacy-preserving only if $|\mathcal{S}_{t}|\geq k_{\min}$; otherwise it must be deferred rather than “fail-open”:
$\mathrm{proceed}(t)=\begin{cases}1,&\text{if }|\mathcal{S}_{t}|\geq k_{\min},\\ 0,&\text{otherwise,}\end{cases}$ (7)
This rule decouples privacy loss from traffic volatility. When DDoS induces long queues at QCaaS and raises $\Delta t_{k}$, aggregation can remain correct by applying the staleness decay $w_{k}(\Delta t_{k})$ while still enforcing (7). The practical failure mode to avoid is a forced reduction of $k_{\min}$ or a switch to plaintext fallback under load; both actions convert an availability incident into a privacy incident. Mitigating this involves mainly three layers:
  1. 1.
    Capacity and scheduling: Implement redundant aggregators with any-cast, regulate client access via rate limits and quotas, randomize PQC job timing to prevent side-channels, and prioritize ongoing rounds to preserve $|\mathcal{S}_{t}|$.
  2. 2.
    Transport and anonymity: Treat relays as honest-but-curious, use anonymous publication and PIR-style access [khan2025controlledpub, khan2022qapir], manage high-contention with quantum collision detection to thwart adversarial actions [khan2021collision], and delegate quantum tasks with concealed computation to protect compute nodes’ knowledge [zaman2023concealed].
  3. 3.
    Content reduction and privacy budgets: Limit symbol exposure by using meaning-level representations [khalid2023qsemantic], hold rounds until $|\mathcal{S}_{t}|\geq k_{\min}$ and apply secure aggregation, and ensure per-round differential privacy to stay within the budget during attacks.
For NISQ deployments, addressing queue sensitivity and scarcity, and integrating these controls with the quantum stack from scratch would be an ideal approach [khalid2024quantumnetwork]. DDoS mainly affects QFL availability but also risks privacy via amplified metadata and timing traces.
Beyond focused attacks such as DDoS, QFL deployments are also exposed to several other distributed threats that can erode confidentiality, integrity, and trust such as :
  • Collusion attacks: Groups of malicious or semi-honest clients can collaborate to reconstruct global model updates, infer private data, or bias aggregation. In classical FL, this already affects differential privacy; in QFL, the risk is higher due to fewer active quantum clients. Coordinated, correlated PQC updates can also deceive anomaly detectors. Mitigation needs secure aggregation protocols that handle collusion and cryptographically shuffle client identities during aggregation [Bonawitz2017SecAgg, khan2025controlledpub].
  • Gradient inference and reconstruction: A curious server or relay may reconstruct training data from gradients. Although quantum noise adds variability, hybrid gradients with classical parts can still leak. Mitigating this involves adding differential privacy noise to both quantum and classical gradients before transmission [mcmahan2017communication], along with secure aggregation to obscure individual contributions [Bonawitz2017SecAgg].
  • Compromise of relay or scheduler nodes: Intermediate nodes managing PQC jobs can tamper with updates, leak metadata, or reorder traffic. Since QFL uses third-party QCaaS or shared infrastructure, these nodes are vulnerable to attacks. Using concealed quantum telecomputation [zaman2023concealed] and anonymous publication mechanisms [khan2025controlledpub, khan2022qapir] can protect client identity and metadata, preventing attribution or targeted delays.
  • Model drift through stale or low-quality updates: In asynchronous QFL, stragglers may submit stale updates using outdated global parameters. If unweighted, these updates can detract the global model from the optimum, reducing accuracy or enabling stealth poisoning attacks. This is worsened on NISQ devices due to calibration drift. Techniques like staleness-weighted aggregation [Xie2019FedAsync] and validation-based reputation scoring can mitigate harmful updates without excluding honest but slow clients.
$\bullet$ Resilience against model and data poisoning: A central question is whether the integration of quantum resources alters the resilience of QFL against widely-known threats such as model and data poisoning. In the classical setting, model poisoning attacks involve adversarial clients submitting deliberately corrupted gradients or updates with the goal of steering global convergence toward an incorrect or backdoored solution. In QFL, a similar threat manifests through the manipulation of variational parameters within PQCs or the injection of malicious amplitude encodings. Although the no-cloning theorem prevents adversaries from perfectly replicating quantum states to amplify their influence, the aggregation rule remains vulnerable to skewed updates [9951322]. Thus, QFL does not eliminate the fundamental risk of model poisoning, but rather shifts its locus from tensor weights to quantum parameters.
Overall, QFL inherits the poisoning vulnerabilities of classical FL while introducing additional complications due to limited error-correction capacity and the opacity of quantum encodings. Yet it also offers distinctive points of resilience: (i) the impossibility of exact state replication restricts large-scale poisoning amplification, and (ii) cryptographic overlays combined with quantum verification techniques can enforce stronger guarantees of update integrity. Consequently, the resilience of QFL against poisoning is not absolute but conditional, hinging on careful protocol design and the alignment of aggregation rules with the constraints of NISQ devices. A comparative overview of vulnerabilities in classical FL versus QFL is summarized in Table VI.
TABLE VI: Comparison of vulnerabilities of FL and QFL against various attacks. QFL tends to be more ($\uparrow$) or similar ($\approx$) vulnerable than classical FL.
Risk class QFL vs FL In QFL: Implementation; Challenges Mitigation Strategy for QFL
Model poisoning $\uparrow$ Manipulation of PQC parameters and encoder calibrations can bias global updates; NISQ noise may mask anomalous drifts, complicating detection [khalid2024quantumnetwork]. Robust/Byzantine aggregation; PQC-parameter anomaly detection; hybrid-key secure aggregation [Blanchard2017Krum, Yin2018Byz, Bonawitz2017SecAgg]
Backdoor inject $\approx$ Gate-parameter steering and trigger states embedded in variational circuits; deeper NISQ circuits increase susceptibility without full QEC [khalid2024quantumnetwork]. Trigger sanitization; parameter attestation for PQC layers; cross-round consistency checks. [zaman2023concealed]
Data poisoning $\approx$ Preparation of corrupted quantum states or adversarial superpositions; heterogeneous noise obscures intent, degrading screening [khalid2024quantumnetwork]. Outlier clipping; robust aggregation; input-state audits / spot verification of measurement statistics. [Yin2018Byz, Blanchard2017Krum, khalid2024quantumnetwork]
Eavesdropping $\uparrow$ Extra hops to QCaaS and relays expose timing, routing, and job metadata; intermediate nodes may be honest-but-curious [zaman2023concealed, khan2025controlledpub, khan2022qapir, khan2021collision]. Concealed telecomputation; anonymous publication / PIR; authenticated channels; relay anonymity. [khan2025controlledpub, khan2021collision]
Availability (DDoS) $\uparrow$ Scarce QPU capacity and long queues make training sensitive to traffic floods; reduced active cohort threatens secure-aggregation thresholds [khalid2024quantumnetwork]. Redundant aggregators; admission control and rate limits; hold-until-$k_{\min}$ policy; staleness-aware async updates. [khalid2024quantumnetwork, Bonawitz2017SecAgg]
@@
Quantum‐specific error mitigation techniques further strengthen security by masking subtle parameter perturbations. Techniques like zero‐noise extrapolation or probabilistic error cancellation [Cerezo2021] can be applied selectively to aggregated model parameters, reducing the impact of malicious noise injections while preserving the integrity of genuine updates. From a cryptographic standpoint, integrating verifiable blind QC protocols [QFL-BlindC] allows the server to challenge clients with hidden circuit specifications, guaranteeing that parameterized unitaries are executed as intended without revealing the model’s secrets. Together, these measures: robust aggregation, fidelity‐based auditing, error‐mitigation filtering, and verifiable execution, offer a layered defense tailored to the hybrid vulnerabilities of QFL.
@@
An assessment of $\mathrm{QUE}$ for secure‐aggregation and attack‐resilient QFL schemes shows uniformly high occupancy due to their use of shallow embedding circuits. For example, the Paillier‐based asynchronous framework in [QFL-AsynchAgg] runs on a minimal two‐qubit, two‐layer ansatz ($\mathrm{Q}_{alloc}=2$, $L=2$), yielding $\mathrm{QUE}=1.0$ and matching its lock‐free aggregation strategy. Post‐quantum MPC in [QFL-SecAggPostQ1] and lattice‐masked secret sharing in [QFL-SecAggPostQ2] both employ four‐qubit, three‐layer circuits ($\mathrm{Q}_{alloc}=4$, $L=3$) with $\mathrm{QUE}=1.0$, justifying synchronous updates to amortize cryptographic overhead. The blockchain‐anchored protocol in [gharavi2025pqbflpostquantumblockchainbasedprotocol] uses a three‐qubit, two‐layer design ($\mathrm{QUE}=1.0$) and also opts for synchronous aggregation to ensure ledger consistency. The adversarial defense in [maouaki2025qfalquantumfederatedadversarial] leverages a four‐qubit, four‐layer QNN ($\mathrm{QUE}=1.0$) and employs asynchronous FedAvg with local perturbation to balance robustness and throughput. Despite their strong QUE, these protocols share key limitations. These models generally ignore client dropout and dynamic participation, lack adaptive privacy‐budget tuning, and incur substantial computational and communication costs from heavy cryptographic primitives. Future work should explore lightweight quantum‐safe primitives, robust dropout‐tolerant aggregation, and real‐time privacy budget adjustment to enhance scalability and resilience in practical QFL deployments.